# largesymbols"0D largesymbols"0D

### v3

@@ 3.2.2 The parameter prior
When generating the training datasets, we draw $\boldsymbol{\theta}^{(1)}$, $\boldsymbol{\theta}^{(2)}$, …, $\boldsymbol{\theta}^{(L)}$ from a distribution of $\boldsymbol{\theta}$. As will be explained in Section 3.3, this distribution of $\boldsymbol{\theta}$ effectively plays the role of a prior for $\boldsymbol{\theta}$. We want the prior to be sufficiently broad to cover possible values of $\boldsymbol{\theta}$ in real applications, but not too broad so that many training datasets become unrealistic and thus wasteful. Below, we explain the prior used in our pretrained NNE. We have also tried several alternative priors. The pretrained NNE’s estimates in applications (Section 4) are not sensitive to the prior specification.
Unlike Bayesian estimation on a single dataset, a unique challenge in specifying our prior is that we need to accommodate datasets with varying dimensions ($d^{(prod)}$, $d^{(cons)}$, $d^{(ads)}$). Consider key statistics such as buy rate and search rate. In real applications, there is no inherent reason for us to expect these key statistics to vary systematically across datasets of different dimensions. Therefore, among training datasets, we would like the key statistics not to vary systematically by data dimensions, either.
This desire rules out common priors such as $\beta_{k}\sim N(0,1)$ for $k=1,...,d^{(prod)}$. With this prior, the buy rates would be systematically higher for simulated datasets with larger $d^{(prod)}$. To see why, note this prior implies $\lVert\boldsymbol{\beta}\rVert^{2}\sim\chi_{d^{(prod)}}^{2}$ which has an expectation equal to $d^{(prod)}$. Consider a simple case where $\boldsymbol{x}_{ij}^{(prod)}\sim N(\boldsymbol{0},\boldsymbol{I})$. Then, the variance of $\boldsymbol{\beta}^{\prime}\boldsymbol{x}_{ij}^{(prod)}$ for any given $\boldsymbol{\beta}$ equals $\lVert\boldsymbol{\beta}\rVert^{2}$, which is expected to increase with $d^{(prod)}$. A larger variance makes it more likely for at least one product to be better than outside option, leading to a higher buy rate.
@@
The prior (2) can also be understood as letting the total impact of product attributes (measured by $\lVert\boldsymbol{\beta}\rVert$; recall that all attributes are standardized) be invariant to $d^{(prod)}$. This means that when there are more product attributes, we expect that each individual attribute has a lower chance to have a large effect on utility. This expectation seems to align with what we see in real applications, where typically only a few “key” product attributes have substantial effects on consumers’ search and purchase decisions.
The discussion so far has left out the priors for the intercepts $\alpha_{0}$ and $\eta_{0}$, which directly affect search cost and outside utility. We specify the priors of these intercepts as normally distributed. Their means and variances are specified to generate realistic search rates and buy rates in training datasets (neither too small nor too close to 1).
@@ 3.2.3 Attributes ($\boldsymbol{x}$)
To generate a training dataset ${\cal D}^{(\ell)}$, we first draw $\boldsymbol{x}^{(\ell)}$ from a distribution of $\boldsymbol{x}$. Next, given $\boldsymbol{x}^{(\ell)}$ and $\boldsymbol{\theta}^{(\ell)}$, we use the search model to simulate $\boldsymbol{y}^{(\ell)}$. Finally, we let ${\cal D}^{(\ell)}=\{\boldsymbol{x}^{(\ell)},\boldsymbol{y}^{(\ell)}\}$. This procedure requires us to specify a distribution of $\boldsymbol{x}$. Conceptually, this distribution should be broad enough to cover common values of $\boldsymbol{x}$ in real applications, but not too broad so that many training datasets become unrealistic and thus wasteful. Below, we describe the distribution of $\boldsymbol{x}$ currently used by us. It is important to note that the pretrained NNE’s estimates in applications (Section 4) are not sensitive to how we specify this distribution.
First, we draw the size of $\boldsymbol{x}$. The size is determined by $d^{(prod)}$, $d^{(ads)}$, $d^{(cons)}$, $J$, and $n$. All these numbers are picked randomly from their respective ranges to which we would like the pretrained NNE to be applicable. Our current choices of these ranges are given in Table 1 (in Section 2.2). A special note is for $n$. We cap $n$ in pretraining (currently at 2e4), because we cannot simulate a dataset of infinite size. However, this cap does not restrict the applications of pretrained NNE, because we can always split a larger dataset and average the estimates across the splits.
Next, we draw the attributes in $\boldsymbol{x}$. There are three features of $\boldsymbol{x}$ commonly seen in real applications that we would like to incorporate. The first feature is that many attributes show substantial consumer-level variations. For example, some consumers see prices systematically higher than other consumers. Therefore, we draw $\boldsymbol{x}$ with two components: a within-consumer component plus a cross-consumer component (the latter is also known as consumer fixed effects). An exception here is $\boldsymbol{x}_{i}^{(cons)}$, which by definition only has the cross-consumer component.
@@
3.3 Link to Bayesian estimation
The eventual test for a pretrained model is to see how it performs in applications. We do so for the pretrained NNE in Section 4 and find that it provides accurate and sensible estimates. That said, from a theoretical standpoint, it is useful to understand the econometric meaning of the estimate by pretrained NNE. Below, we will explain that pretrained NNE has a familiar Bayesian interpretation, in that it calculates the posterior mean of $\boldsymbol{\theta}$ given $\boldsymbol{m}$. In words, it provides the best estimate of the structural model’s parameter $\boldsymbol{\theta}$ given the data patterns $\boldsymbol{m}$.
Continuing with the notation in (1), let ${\cal D}=\{\boldsymbol{x},\boldsymbol{y}\}$ denote a dataset. As before, $\boldsymbol{x}$ and $\boldsymbol{y}$ here collect the observations across all consumers. The vector $\boldsymbol{m}=\boldsymbol{h}({\cal D})$ denotes a set of data patterns. We start with a result with the standard (non-pretrained) NNE, studied in Wei and Jiang (2024). The main difference between standard NNE and pretrained NNE is that the former is trained for a specific real dataset. The training examples do not simulate $\boldsymbol{x}$ but directly use the observed $\boldsymbol{x}$ from that real dataset.
The standard NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m},\boldsymbol{x})$ as $L\rightarrow\infty$ (see Wei and Jiang 2024). Note that the result does not require large $n$ and holds for any value of $n$. Also note that we have made it explicit that the expectation conditions on a specific $\boldsymbol{x}$ [^4] The notations of Wei and Jiang (2024) omit the conditioning on $\boldsymbol{x}$ because the paper assumes an underlying probability space that is already conditional on the observed $\boldsymbol{x}$. The $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m},\boldsymbol{x})$ is known as a limited-information Bayesian posterior mean (e.g., Kim 2002), where “limited-information” indicates that the posterior relies on less information than the full data. To compare, the full-information posterior is $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{y},\boldsymbol{x})$. Ideally, we would like to compute the full-information posterior, but when doing so is difficult, the limited-information posterior provides a practical alternative.
We now turn to the pretrained NNE. In this case, the training examples do not assume a specific $\boldsymbol{x}$, but instead simulate $\boldsymbol{x}$ from a distribution (as described in Section 3.2). Consequently, the estimate by the pretrained NNE converges to $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$ as $L\rightarrow\infty$. Importantly, note that the expectation no longer conditions on $\boldsymbol{x}$.
Given this convergence result, a natural question is how to interpret $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$. Formally, it is another limited-information posterior mean. For a more intuitive interpretation, consider a practical scenario where a researcher wants to apply the search model to a company’s data. For privacy reasons, the company provides the researcher with only aggregate data patterns $\boldsymbol{m}$, instead of the full data $\{\boldsymbol{x},\boldsymbol{y}\}$. In this case, our best estimate for the search model parameter will be precisely $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$. By definition,
[…]
In equation (3), the term $P(\boldsymbol{m}|\boldsymbol{\theta},\boldsymbol{x})$ is defined by the search model, i.e., the likelihood of observing the data patterns $\boldsymbol{m}$ given $\boldsymbol{\theta}$ and $\boldsymbol{x}$. The term $P(\boldsymbol{\theta},\boldsymbol{x})$ must be specified by the researcher as a prior. Typically, the prior for $\boldsymbol{\theta}$ is specified independently of $\boldsymbol{x}$: $P(\boldsymbol{\theta},\boldsymbol{x})=P(\boldsymbol{\theta})P(\boldsymbol{x})$. These two components of the prior, $P(\boldsymbol{\theta})$ and $P(\boldsymbol{x})$, correspond to the distributions we use to draw $\boldsymbol{\theta}$ and $\boldsymbol{x}$ during the pretraining (described in Section 3.2).
This link to Bayesian estimation guides our specification of $\boldsymbol{m}$. Because the pretrained NNE computes $\mathbb{E}(\boldsymbol{\theta}|\boldsymbol{m})$, $\boldsymbol{m}$ should include the data patterns that are informative about $\boldsymbol{\theta}$. Meanwhile, we would like the computation of $\boldsymbol{m}$ to be inexpensive. These considerations have led us to use the coefficients from various reduced-form regressions of $\boldsymbol{y}$ on $\boldsymbol{x}$ (see Section 3.1).
@@ 4 Applications
We evaluate the pretrained NNE’s performance with 12 real datasets. In particular, we compare it to SMLE, which is the prevailing approach to estimate search models. Below, we first briefly describe each dataset. Then, we conduct Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices. Finally, we estimate the search model on the real data entirely.
Before proceeding, we briefly describe existing approaches to estimate search models, for readers unfamiliar with the literature. Standard SMLE with accept-reject simulations is infeasible for search models because it requires prohibitive computational costs. A common workaround in literature is to smooth the simulated likelihood function (e.g., Honka and Chintagunta 2016, Ursu 2018). However, estimates are sensitive to smoothing factors and tuning these factors requires costly Monte Carlo experiments. More recently, Jiang et al. (2021) and Chung et al. (2023) extend the Geweke-Hajivassiliou-Keane (GHK) simulation algorithm to the likelihood of search models. The GHK-based SMLE does not require smoothing, but is much more complex to implement. We will consider both the smoothed SMLE and the GHK-based SMLE.
@@ 4.1 Datasets
Our 12 datasets come from four different settings (an anonymous online retailer, Expedia, JD, and Trivago). Table 2 provides summary statistics. We see that these datasets vary substantially in the number of search sessions, numbers of attributes, buy rate, and search rate. So they cover a fairly wide range of real-world scenarios.
For ease of presentation, the estimates reported in this section assume that each attribute is mean-zero with a standard deviation of 0.2 (instead of 1). The 0.2 standard deviation makes coefficients ($\boldsymbol{\beta}$, $\boldsymbol{\eta}$, $\boldsymbol{\alpha}$) larger and thus more visible in figures. For estimation, all attributes are standardized to unit variance before being fed into the pretrained NNE (see Section 2.2).
@@
Our first dataset consists of consumer searches and purchases in an anonymous large online store, provided by the marketing company REES46. The data is available on Kaggle.com. [^5] https://www.kaggle.com/datasets/mkechinov/ecommerce-behavior-data-from-multi-category-store The store sells products across many categories. This dataset focuses on laptops. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and five brand dummies: Asus, HP, Lenovo, MSI, and Xiaomi (the baseline brand is Acer). There are no consumer or advertising attributes. There are a very large number of search sessions. To ease the computational burden of SMLE, we take a subsample of 10,000 search sessions.
@@
The International Conference on Data Mining (ICDM 2013) organized a contest to predict consumer searches and bookings for hotels on Expedia. The data for the contest is publicly available on Kaggle.com. [^6] https://www.kaggle.com/competitions/expedia-personalized-sort/overview It has been used in several papers to study consumer search behaviors (Ursu 2018, Greminger 2022, Compiani et al. 2023). The data includes multiple travel destinations. “Destination 1” refers to the largest destination. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include star rating, review score, a chain indicator, location score, price, and promotion. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The advertising attribute $\boldsymbol{x}_{ij}^{(ads)}$ includes the ranking of the hotel.
@@
This dataset consists of the search sessions for the second largest destination in the Expedia data from ICDM 2013. There are only $n=897$ search sessions, slightly below the minimum threshold of $n$ set for the pretrained NNE (see Table 1). Nonetheless, we decide to include the dataset, as it allows us to observe how the pretrained NNE performs slightly outside its intended range of applications.
@@
The Analytics at Wharton Initiative also provides data on hotel searches and bookings on Expedia. [^7] https://analytics.wharton.upenn.edu/open-research-opportunities/predicting-optimizing-consumer-response-to-product-search-results/ The data is not openly available but can be accessed after a request with the Initiative. The data includes four travel destinations (Cancun, Manhattan, Paris and Budapest). Cancun is the largest destination.
@@
This dataset consists of the search sessions for the second largest destination (Manhattan) in the Expedia data from Analytics at Wharton.
@@
The 2020 MSOM Data Driven Research Challenge provided data on consumer experiences on JD.com. The data was made available to INFORMS members in 2019.888https://connect.informs.org/msom/events/datadriven2020 JD.com is China’s second largest business-to-consumer online retail platform. The data covers many aspects of consumer experiences on the platform, from website browsing to order fulfillment (see Shen et al. 2020 for details). We focus on consumer search and purchase activities. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price, a dummy for first-party sellers, and two unnamed attributes. There are no advertising attributes.
The JD.com data includes three channels: desktop, mobile, and app. This dataset focuses on the desktop channel.
@@
The ACM RecSys Challenge 2019 provides consumer data on Trivago.com. [^9] https://recsys.acm.org/recsys19/challenge/ The data is not standard for search models because purchases are not very well defined in this context, as explained below. Nonetheless, we decide to include the data, because it allows us to observe how the pretrained NNE performs somewhat outside the intended context of the search model.
More specifically, Trivago is a hotel meta-search engine that aggregates information from hotels and travel websites (similar to how Google Flights aggregates flight information). On Trivago.com, consumers receive a list of hotels. They can click on any hotel to see more information. They can also click a “View Deal” button that directs them to the external site for possible booking. But what happens at external sites is not observed in the data. We treat a click for more information as a search and a “View Deal” click as a purchase. Sometimes a consumer clicks “View Deal” for multiple hotels, in which case we randomly pick one as the purchase. Cases where consumers click “View Deal” without first clicking for more information are dropped (about 5%).
The Trivago data includes search sessions from both the desktop channel and mobile channel. This dataset focuses on the desktop channel.

### v4

@@ 3.3.2 The parameter prior
When generating the training datasets, we draw $\boldsymbol{\theta}^{(1)}$, $\boldsymbol{\theta}^{(2)}$, …, $\boldsymbol{\theta}^{(L)}$ from a distribution of $\boldsymbol{\theta}$. As explained in Section 3.2, this distribution of $\boldsymbol{\theta}$ effectively plays the role of a prior for $\boldsymbol{\theta}$. We want the prior to be sufficiently broad to cover possible values of $\boldsymbol{\theta}$ in real applications, but not so broad that many training datasets become unrealistic and thus wasteful. Below, we explain the prior used in our pretrained NNE. We have also tried several alternative priors. The pretrained NNE’s estimates in applications (Section 4) are not sensitive to the prior specification.
Unlike Bayesian estimation on a single dataset, a unique challenge in specifying our prior is that we need to accommodate datasets with varying dimensions ($d^{(prod)}$, $d^{(cons)}$, $d^{(ads)}$). Consider key statistics such as buy rate and search rate. In real applications, there is no inherent reason for us to expect these key statistics to vary systematically across datasets of different dimensions. Therefore, across training datasets, we would like the key statistics not to vary systematically by data dimensions, either.
This desire rules out common priors such as $\beta_{k}\sim N(0,1)$ for $k=1,...,d^{(prod)}$. With this prior, the buy rates would be systematically higher in training datasets with larger $d^{(prod)}$. To see why, note this prior implies $\lVert\boldsymbol{\beta}\rVert^{2}\sim\chi_{d^{(prod)}}^{2}$ which has an expectation equal to $d^{(prod)}$. Consider a simple case where $\boldsymbol{x}_{ij}^{(prod)}\sim N(\boldsymbol{0},\boldsymbol{I})$. Then, the variance of $\boldsymbol{\beta}^{\prime}\boldsymbol{x}_{ij}^{(prod)}$ for any given $\boldsymbol{\beta}$ equals $\lVert\boldsymbol{\beta}\rVert^{2}$, which in expectation increases with $d^{(prod)}$. A larger variance makes it more likely for at least one product to be better than outside option, leading to a higher buy rate.
@@
Intuitively, the prior (3) can also be understood as letting the overall impact of product attributes (as measured by $\lVert\boldsymbol{\beta}\rVert$) be invariant to $d^{(prod)}$. This means that when there are more product attributes, we expect that each individual attribute has a lower chance to have a large effect on utility. This belief seems to align with what we see in real applications, where typically only a few “key” product attributes have substantial effects on consumers’ search and purchase decisions.
The discussion so far has left out the priors for the intercepts, $\alpha_{0}$ and $\eta_{0}$, which directly affect search cost and outside utility. We use normal priors for these intercepts, with their means and variances specified to produce realistic search rates and buy rates in training datasets (neither too small nor too close to 1).
@@ 3.3.3 Attributes ($\boldsymbol{x}$)
The generation of training examples requires us to specify a distribution of $\boldsymbol{x}$. Specifically, to generate a training dataset ${\cal D}^{(\ell)}$, we first draw $\boldsymbol{x}^{(\ell)}$ from this distribution. Then, given $\boldsymbol{x}^{(\ell)}$ and $\boldsymbol{\theta}^{(\ell)}$, we use the search model to simulate $\boldsymbol{y}^{(\ell)}$, and let ${\cal D}^{(\ell)}=\{\boldsymbol{x}^{(\ell)},\boldsymbol{y}^{(\ell)}\}$. Conceptually, the distribution of $\boldsymbol{x}$ should be broad enough to cover common values of $\boldsymbol{x}$ in real applications, but not so broad that many training datasets become unrealistic and thus wasteful. Below, we describe the distribution of $\boldsymbol{x}$ currently used by us. It is important to note that the pretrained NNE’s estimates in applications (Section 4) are not sensitive to the specification of this distribution.
First, we draw the size of $\boldsymbol{x}$. The size is determined by $d^{(prod)}$, $d^{(ads)}$, $d^{(cons)}$, $J$, and $n$. These numbers are picked at random from their respective ranges to which we would like the pretrained NNE to be applicable. Our current choices of these ranges are given in Table 1 (in Section 2.2). In particular, the upper bound of $n$ in pretraining is currently set at 2e4. However, this cap does not restrict the applications of pretrained NNE, because we can always split a larger dataset and average the estimates across the splits (also discussed in Section 2.2).
Next, we draw the attributes in $\boldsymbol{x}$. In this regard, we incorporate several features of $\boldsymbol{x}$ commonly seen in real applications. The first feature is that many attributes show substantial consumer-level variations. For example, some consumers see prices systematically higher than other consumers. Therefore, we draw $\boldsymbol{x}$ with two components: a within-consumer component plus a cross-consumer component (the latter is also known as consumer fixed effects). The within-consumer component is ignored for $\boldsymbol{x}_{i}^{(cons)}$, which by definition varies only cross consumers.
@@ 4 Applications
We evaluate the pretrained NNE’s performance with 12 real datasets. First, we briefly describe each dataset. Then, we conduct Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices. Finally, we estimate the search model on the real data entirely.
Before proceeding, we briefly describe the existing approaches to estimate search models, for readers unfamiliar with the literature. Standard SMLE with accept-reject simulations is infeasible for search models because it requires prohibitive computational costs. A common workaround in literature is to smooth the simulated likelihood function (e.g., Honka and Chintagunta 2016, Ursu 2018). However, estimates are sensitive to smoothing factors and tuning these factors requires costly Monte Carlo experiments. More recently, Jiang et al. (2021) and Chung et al. (2023) extend the Geweke-Hajivassiliou-Keane (GHK) simulation algorithm to the likelihood of search models. The GHK-based SMLE does not require smoothing, but is much more complex to implement. In this section, we implement both the smoothed SMLE and the GHK-based SMLE.
@@ 4.1 Datasets
Our 12 datasets come from four different sources (an anonymous online retailer, Expedia, JD, and Trivago). Table 2 provides summary statistics. We see that these datasets vary substantially in the number of search sessions, numbers of attributes, buy rate, and search rate. So they cover a fairly wide range of real-world scenarios.
For ease of presentation, the estimates reported in this section assume that each attribute is mean-zero with a standard deviation of 0.2. The 0.2 standard deviation (instead of 1) makes coefficients ($\boldsymbol{\beta}$, $\boldsymbol{\eta}$, $\boldsymbol{\alpha}$) larger and thus more visible in figures. For estimation, all attributes are standardized to unit variance before being fed into the pretrained NNE (see Section 2.2).
@@
Our first data source consists of consumer searches and purchases in an anonymous large online store, provided by the marketing platform REES46. The data are available on Kaggle.com and include many product categories. [^5] https://www.kaggle.com/datasets/mkechinov/ecommerce-behavior-data-from-multi-category-store We focus on laptops in this dataset. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and five brand dummies: Asus, HP, Lenovo, MSI, and Xiaomi (the baseline brand is Acer). There are no consumer or advertising attributes. There are a very large number of search sessions. To ease the computational burden of SMLE, we use a subsample of 10,000 search sessions.
@@
The International Conference on Data Mining (ICDM 2013) organized a contest to predict consumer searches and bookings for hotels on Expedia. The data for the contest are publicly available on Kaggle.com. [^6] https://www.kaggle.com/competitions/expedia-personalized-sort/overview Several papers have used the data to study consumer search behaviors (Ursu 2018, Greminger 2022, Compiani et al. 2023). The data include multiple travel destinations. “Destination 1” refers to the largest destination. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include star rating, review score, a chain indicator, location score, price, and promotion. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The advertising attribute $\boldsymbol{x}_{ij}^{(ads)}$ includes the ranking of the hotel.
@@
This dataset consists of the search sessions for the second-largest destination in the Expedia data from ICDM 2013. There are only $n=897$ search sessions, slightly below the lower bound of $n$ set for the pretrained NNE (see Table 1). Nonetheless, we choose to include the dataset, because it allows us to test the pretrained NNE slightly outside its intended range of applications.
@@
The Analytics at Wharton Initiative also provides data on hotel searches and bookings on Expedia. [^7] https://analytics.wharton.upenn.edu/open-research-opportunities/predicting-optimizing-consumer-response-to-product-search-results/ The data are not openly available but can be accessed after a request with the Initiative. The data include four travel destinations (Cancun, Manhattan, Paris and Budapest). Cancun is the largest destination.
@@
This dataset consists of the search sessions for the second-largest destination (Manhattan) in the Expedia data from Analytics at Wharton.
@@
The 2020 MSOM Data Driven Research Challenge provided data on consumer experiences on JD.com. The data were made available to INFORMS members in 2019.888https://connect.informs.org/msom/events/datadriven2020 JD.com is China’s second largest business-to-consumer online retail platform. The data cover many aspects of consumer experiences on the platform, from website browsing to order fulfillment (see Shen et al. 2020 for details). We focus on consumer search and purchase activities. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price, a dummy for first-party sellers, and two unnamed attributes. There are no advertising attributes.
The JD.com data include three channels: desktop, mobile, and app. This dataset focuses on the desktop channel.
@@
The ACM RecSys Challenge 2019 provided consumer browsing data on Trivago.com. [^9] https://recsys.acm.org/recsys19/challenge/ The data are not standard for search models because purchases are not very well defined in this context, as we will explain below. Nonetheless, we choose to include the data because doing so allows us to test the pretrained NNE slightly outside its intended context of applications.
Trivago is a hotel meta-search engine that aggregates information from hotels and travel websites (similar to how Google Flights aggregates flight information). On Trivago.com, consumers receive a list of hotels. They can click on any hotel to see more information. They can also click a “View Deal” button that directs them to an external website for possible booking. But what happens at external websites is not observed in the data. We treat a click for more information as a search and a “View Deal” click as a purchase. Sometimes a consumer clicks “View Deal” for multiple hotels, in which case we randomly pick one as the purchase. Cases where consumers click “View Deal” without first clicking for more information are dropped (about 5%).
The Trivago data include search sessions from both the desktop channel and mobile channel. This dataset focuses on the desktop channel.